# Moving Aggregate Modified Autoregressive Copula-Based Time Series Models (MAGMAR-Copulas)

### v2

@@ 2.3 Copula-based Time Series models with an arbitrary autoregressive order
If no superscript is given, then conditioning on the second variable is implied. The conditional distribution $F_{U_{t}|U_{t-1}}$ may be rewritten as $F_{U_{t}|U_{t-1}}(u_{t},u_{t-1})=C_{\phi}(u_{t}|u_{t-1})=h^{2}_{\phi_{1}}(u_{t},u_{t-1})$. The same applies to $U_{t-2}|U_{t-1}$: $F_{U_{t-2}|U_{t-1}}(u_{t-2},u_{t-1})=C_{\phi}(u_{t-2}|u_{t-1})=h^{1}_{\phi}(u_{t-1},u_{t-2})$. The $h$-functions can be derived from the associated copula. Their calculation amounts to taking the derivative of copulas. Often the $h$-functions can be derived analytically. (Explicit forms of $h$-functions and their inverses can be found in Smith et al. (2010) for the normal, t, clayton, gumbel and galambos copula.) An advantage of using the $h$-functions is that all conditional distributions can be decomposed into a function of $h$-functions. E.g. the distribution of $(U_{t}|U_{t-1},U_{t-2})$ which is needed for the partial distribution of $U_{t}$ and $U_{t-3}$, can be calculated as (Aas et al., 2009, Joe, 2014)
$F_{U_{t}|U_{t-1},U_{t-2}}(u_{t}|u_{t-1},u_{t-2})=h_{\phi_{2}}^{2}\big{(}h^{2}_{\phi_{1}}(u_{t},u_{t-1}),h^{1}_{\phi_{1}}(u_{t-1},u_{t-2})\big{)}.$
The notion of decomposing a conditional distribution into a function of $h$-functions is named Rosenblatt (forward/backward) function (McNeil and Bladt, 2022). The decomposition of $U_{t}|U_{t-1},\ldots,U_{t-p}$ can be called Rosenblatt backward and the decomposition of $U_{t-p}|U_{t-p+1},\ldots,U_{t}$ Rosenblatt forward function (inspired by the Rosenblatt transform from Rosenblatt (1952)). For $p=1$ the Rosenblatt functions are just the corresponding $h$-functions, $R_{1}^{1}=h_{1}^{1}$, $R_{1}^{2}=h_{1}^{2}$. McNeil and Bladt (2022) define the Rosenblatt forward and backward functions for $p>1$, $R_{p}:(0,1)^{p}\times(0,1)\rightarrow(0,1)$, recursively as follows:
@@
$R_{p}^{2}(\mathbf{u},x)=$ $h_{p}^{2}\left(R_{p-1}^{2}(\mathbf{u}_{-p},x),R_{p-1}^{1}(\mathbf{u}_{-p},u_{p})\right).$ (8)
Here $\mathbf{u}_{-j}$ denotes $\mathbf{u}_{-j}=(u_{1},\ldots,u_{j-1},u_{j+1},\ldots,u_{k})$. The second argument of the Rosenblatt functions, $x$, is the actual variable that the Rosenblatt functions are evaluated at. The conditioning variables are contained in the vector $\mathbf{u}$. Using the newly introduced notation, the D-vine copula density (Eq. 6) can be rewritten as (McNeil and Bladt, 2022),
@@
$U_{t}=R^{-1}_{p}((U_{t-p},\ldots,U_{t-1}),w_{t}),\ w_{t}\stackrel{{\scriptstyle iid}}{{\sim}}U(0,1).$ (9)
Here $R_{p}$ is the Rosenblatt function corresponding to the $(p+1)$-dimensional copula $C_{\phi}$ from Eq. 5. In this notation, the Markov($p$) restriction becomes obvious. The model can also be understood as a generalized AR($p$)-model. In this case, however, the function connecting the past variables, $U_{t-1},\ldots,U_{t-p}$, innovations $w_{t}$ and the current observations $U_{t}$ is not necessarily linear. For the case where $p=1$, the updating equation becomes
$U_{t}=h_{\phi}^{-1}(w_{t},U_{t-1}),\ w_{t}\stackrel{{\scriptstyle iid}}{{\sim}}U(0,1).$ (10)
Here $h_{\phi}$ is the $h$-function associated to the 2-dimensional copula $C_{\phi}$ from Eq. 4. Next, we are going to discuss the extension of McNeil and Bladt (2022) to allow for infinite partial dependence.
@@ 2.4 Review of the extension by McNeil & Bladt
McNeil and Bladt (2022) allow for infinite partial dependence in copula-based time series models by letting the autoregressive order, $p$, be infinite and reparametrizing the joint density with a finite amount of parameters. The parameters are chosen in such a way that they represent a classical ARMA($p,q$) process with finite $p$ and $q$. To model a time series, the following steps are employed. First, the autoregressive and the moving average order, $p$ and $q$, respectively, are determined by minimizing model selection criteria (e.g. AIC). Next, the partial autocorrelations of the chosen ARMA($p,q$) process are parametrized in terms of the autoregressive and the moving average parameters, $\phi_{1},\ldots,\phi_{p},\theta_{1},\ldots,\theta_{q}$. Using the one-to-one relation between autocorrelations and Kendall’s $\tau$, the Kendall’s partial autocorrelations are obtained. In a third step, the partial copulas in Eq. 6 are reparametrized by Kendall’s $\tau$ and consecutively by the AR and MA parameters, which can be estimated by maximum likelihood.
One disadvantage of their extension is that the Kendall’s $\tau$ reparametrization only works for one-parametric copulas, since only for one-parametric copulas, there is a one-to-one relation. This excludes, for example, the t-copula, which is of particular relevance in practice (see e.g. Loaiza-Maya et al., 2018, Pappert and Arsova, 2023). Reparametrization for copulas with more than one parameter is more cumbersome.
@@ 2.5 Review of the extension by Joe
Now we want to briefly introduce the model of Joe (2014), which is to the best of our knowledge the earliest copula extension of the ARMA model. The model is quite similar to the MAGMAR-copula model, although it is not the same. The model with uniform marginals is given by the following updating equations:
@@
where $\varepsilon_{t}\stackrel{{\scriptstyle iid}}{{\sim}}U(0,1)$ are the common innovations, $h$ is the conditional quantile function of a $(p+1)$-variate copula and $g$ is the conditional quantile function of a $(q+1)$-variate copula. It is required that the copula corresponding to $h$ is translationally invariant and that the $q$-variate marginal distribution of the copula corresponding to $g$ is the independence copula. The idea is inspired by the Poisson ARMA process of McKenzie (1988): First create a process corresponding to an AR process, Eq. 11, and then feed this process to an MA-like filter together with the common innovations, Eq. 12. In our proposed MAGMAR-copula model, we too attempt to mix AR-type and MA-type copula processes. The difference, however, is that we feed the MA-type process (what we call ’MAG’) into the AR-type filter. In other words, in the copula extension to ARMA models by Joe (2014) the latent process is an AR-type process whereas in the MAGMAR-copula model, the latent process is an MA-like model.
@@ 3 MAGMAR-Copulas
In this section, we introduce and discuss the moving aggregate modification to copula-based time series models. For illustration, the autoregressive and moving aggregate orders are set to one. Results for arbitrary orders $p$ and $q$ can be derived in a similar fashion; Some results with arbitrary orders are given in the Appendix.
@@ 3.1 Introduction to the model
Our proposed approach enabling the model to overcome the Markov restriction amounts to modifying the autoregressive updating equation, Eq. 10, by a moving aggregate part. The moving aggregate part is, similar to classical, linear ARMA processes, a function of the current and past innovations. It is given as an inverse conditional copula, which we parametrize with $\theta$. So instead of $w_{t}$ in Eq. 10, we have $h_{\theta}^{-1}(w_{t},w_{t-1})$. We call the model Moving Aggregate Modified Autoregressive Copula-Based Time Series Model (MAGMAR-Copula model). In the following exposition, we assume that 1) all copulas are differentiable, 2) that they are within the Fréchet-Hoeffding bounds (see Joe, 2014, Chapter 2.9) and 3) that the inverses of $h$-functions associated to all used copulas exist.
Definition 1 (MAGMAR(1,1)11(1,1)( 1 , 1 )-Copula model).
@@
Let $h^{-1}_{\phi}:(0,1)^{2}\rightarrow(0,1)$ and $h_{\theta}^{-1}:(0,1)^{2}\rightarrow(0,1)$ be inverse $h$-functions associated to copulas $C_{\phi}$ and $C_{\theta}$, which are parametrized by parameters $\phi$ and $\theta$, respectively. The time series $\{U_{t}\}_{t\in\mathbb{Z}}$ following the updating equation
$U_{t}=h_{\phi}^{-1}(h_{\theta}^{-1}(w_{t},w_{t-1}),U_{t-1}),\ \ \ w_{t}\stackrel{{\scriptstyle iid}}{{\sim}}U(0,1)$ (13)
@@
is a MAGMAR(1,1)11(1,1)( 1 , 1 )-Copula time series. The corresponding model is the MAGMAR(1,1)11(1,1)( 1 , 1 )-Copula model. The copula $C_{\phi}$ is the AR-Copula, while $C_{\theta}$ is the MAG-Copula. $h_{\theta}^{-1}(w_{t},w_{t-1})$ is the MAG-part.
The $h$-functions associated to the AR and MAG-copula can be derived from the respective copulas using Eq. 7. The family of the copulas (e.g. normal, t, gumbel, etc.) has to be chosen. Parameters $\phi$ and $\theta$ have to be set/estimated. We note that the MAG-part, that is used here, was also derived in McNeil and Bladt (2022) as the ’causal representation’ for the classical copula-based time series models. The new contribution is that we use this representation to construct the new MAGMAR-Copula model.
@@
Before giving the relations to other models and deriving distributional properties, we proceed to discuss implications of the model equation and further explain the model. First, we note that $C_{\theta}$ is not the copula of $w_{t}$ and $w_{t-1}$. The innovations $w_{t}$ and $w_{t-1}$ are independent and their copula is the independence copula, $F_{w_{t},w_{t-1}}=\Pi$. The expression $h_{\theta}(w_{t},w_{t-1})$ is to be understood as a function of two random variables $w_{t}$ and $w_{t-1}$, just as $\varepsilon_{t}-\theta\varepsilon_{t-1}$ in the classical ARMA model is a function of two random variables $\varepsilon_{t}$ and $\varepsilon_{t-1}$. Second, The copula $C_{\phi}$ is also not the copula of $U_{t}$ and $U_{t-1}$, however it is connected to the copula of the two random variables. This is analogous to the classical ARMA(1,1) model, where the partial autocorrelation of $X_{t-1}$ and $X_{t}$ is not solely determined by the AR(1) parameter but also by the MA(1) parameter. (The influence of the MAG-Copula and the AR-Copula to the dependence of $U_{t}$ and $U_{t-1}$ becomes clear when we derive the joint conditional distribution of $U_{t}$ and $U_{t-1}$ given $\mathcal{F}_{t-2}$.) The expression $h_{\phi}(h_{\theta}^{-1}(w_{t},w_{t-1}),U_{t-1})$ may also be understood as a function of random variables. To verify that the MAGMAR(1,1)11(1,1)( 1 , 1 ) model equation actually yields the expected behavior to have no Markov restriction i.e. infinite autoregressive order, the model equation can be iterated. To this aim, an explicit form for the innovations is necessary. It can be directly obtained from Eq. 13 as
@@
Hence, in general, each past observation, $U_{t-i},i=1,2,\ldots$ has a non-independent influence on $U_{t}$. The MAG($\infty$) representation can be derived as
@@
As noted before, there also exists a MAG($\infty$) representation for the case with no moving aggregate extension. The representation was derived in McNeil and Bladt (2022) as ’causal representation’. It is easy to see that for $C_{\theta}=\Pi$, the two representations are equal. This is in line with proposition 2, which proposes that for $C_{\theta}=\Pi$, the MAGMAR(1,1)-Copula model is equal to the classical copula-based time series model. The proposition and the proof are given in Sect. 3.2. In Sect. 3.3.1, it is explicitly shown that the MAG($\infty$) representation of the MAGMAR-Copula time series converges for the case where the AR-copula, $C_{\phi}$, is the normal copula.
Last, we want to introduce the MAGMAR$(p,q)$-Copula model with $p$ autoregressive and $q$ moving aggregate orders. Using the Rosenblatt functions (Eq. 8) It can be formulated as follows:
Definition 2 (MAGMAR$(p,q)$-Copula Model).
Let $R_{\phi_{1},\ldots,\phi_{p}}:(0,1)^{p+1}\rightarrow(0,1)$ be a Rosenblatt function associated to a D-vine copula $C_{\phi_{1},\ldots,\phi_{p}}$, which is parametrized by parameters $\phi_{1},\ldots,\phi_{p}$. Furthermore, let $R^{-1}_{\theta_{1},\ldots,\theta_{q}}:(0,1)^{q+1}\rightarrow(0,1)$ be an inverse Rosenblatt function associated to a D-vine copula $C_{\theta_{1},\ldots,\theta_{q}}$, which is parametrized by parameters $\theta_{1},\ldots,\theta_{q}$. The time series $\{U_{t}\}_{t\in\mathbb{Z}}$ following the updating equation
is a MAGMAR$(p,q)$-time series. The corresponding model is the MAGMAR$(p,q)$-Copula model. The copula $C_{\phi_{1},\ldots,\phi_{p}}$ is the AR-Copula. The copula $C_{\theta_{1},\ldots,\theta_{q}}$ is the MAG-Copula. The LHS of Eq. 16 is the AR-part and the RHS is the MAG-part.
The Rosenblatt functions in Eq. 16 ought to be understood as a composition of $h$-functions (as given in Eq. 8). The AR-Rosenblatt function, $R^{1}_{\phi_{1},\ldots,\phi_{p}}$ is composed of $h$-functions $h_{\phi_{1}},\ldots,h_{\phi_{p}}$ while the MAG-Rosenblatt function is composed of $h$-functions $h_{\theta_{1}},\ldots,h_{\theta_{q}}$. Analogous to to the MAGMAR(1,1)11(1,1)( 1 , 1 ) case, the AR-Rosenblatt function is not to be understood as conditional distribution of $U_{t}$ given $U_{t-1},\ldots,U_{t-p}$, but as a mere function, whose arguments are $U_{t},\ldots,U_{t-p}$. The same holds for the MAG-Rosenblatt function and $w_{t},w_{t-1},\ldots,w_{t-q}$. In the case where the copulas, respectively $h$-functions involved in $R^{1}_{\theta_{1},\ldots,\theta_{t-q}}$ are chosen as the independence copula, the conditional distribution of $U_{t}|U_{t-1},\ldots,U_{t-p}$ is given by $R^{1}_{\phi_{1},\ldots,\phi_{p}}$. The MAG($\infty$)∞ ) and the AR($\infty$) representations of the MAGMAR$(p,q)$-Copula can be derived in a similar fashion to the MAGMAR(1,1)11(1,1)( 1 , 1 )-Copula case. This result is given in Appendix D. Next, we give a conceptual comparison between the MAGMAR-Copula model and the copula-ARMA model by McNeil and Bladt (2022).

### v3

@@ 2.2 Copula-based Time Series with Autoregressive Order of One
If no superscript is given, then conditioning on the second variable is implied. Using this notation, the conditional distribution of e.g. $F_{U_{t}|U_{t-1}}$ may be rewritten as $F_{U_{t}|U_{t-1}}(u_{t},u_{t-1})=C_{\phi}(u_{t}|u_{t-1})=h^{2}_{\phi_{1}}(u_{t},u_{t-1})$. Explicit forms of $h$-functions and their inverses can be found in Smith et al. (2010) for the normal, $t$, Clayton, Gumbel and Galambos copula.) An advantage of using the $h$-functions is that all conditional distributions can be decomposed into a composition of $h$-functions, i.e. Rosenblatt (forward/backward) functions (McNeil and Bladt, 2022). The decomposition of $U_{t}|U_{t-1},\ldots,U_{t-p}$ can be called Rosenblatt backward and the decomposition of $U_{t-p}|U_{t-p+1},\ldots,U_{t}$ Rosenblatt forward function (inspired by the Rosenblatt transform from Rosenblatt (1952)). For $p=1$ the Rosenblatt functions are just the corresponding $h$-functions, $R_{1}^{1}=h_{1}^{1}$, $R_{1}^{2}=h_{1}^{2}$. McNeil and Bladt (2022) define the Rosenblatt forward and backward functions for $p>1$, $R_{p}:(0,1)^{p}\times(0,1)\rightarrow(0,1)$, recursively as follows:
@@
$R_{p}^{2}(\mathbf{u},x)=$ $h_{p}^{2}\left(R_{p-1}^{2}(\mathbf{u}_{-p},x),R_{p-1}^{1}(\mathbf{u}_{-p},u_{p})\right),$ (7)
where $\mathbf{u}_{-j}=(u_{1},\ldots,u_{j-1},u_{j+1},\ldots,u_{k})$. The second argument of the Rosenblatt functions, $x$, is the variable that the Rosenblatt functions are evaluated at. The conditioning variables are contained in the vector $\mathbf{u}$. Using the newly introduced notation, the D-vine copula density (Eq. 5) can be rewritten as (McNeil and Bladt, 2022),
@@
$U_{t}=R^{-1}_{p}((U_{t-p},\ldots,U_{t-1}),w_{t}),\quad w_{t}\stackrel{{\scriptstyle iid}}{{\sim}}U(0,1).$ (8)
Here $R_{p}^{-1}$ is the inverse Rosenblatt function corresponding to the $(p+1)$-dimensional copula $C_{\phi}$ from Eq. 4. The inversion is w.r.t. the last argument. Formally, $R_{p}^{-1}(\mathbf{u},x)$ is the solution to $R_{p}(\mathbf{u},x)=x$. In this notation, the Markov($p$) restriction becomes obvious. Next, we are going to discuss the extension of McNeil and Bladt (2022) and Joe (2014) to allow for infinite partial dependence.
@@ 2.3 Review of the extension by McNeil & Bladt
McNeil and Bladt (2022) allow for infinite partial dependence in copula-based time series models by letting the autoregressive order, $p$, be infinite and reparametrizing the joint density with a finite amount of parameters. The parameters are chosen in such a way that they represent a classical ARMA($r,s$) process with finite $r$ and $s$. To model a time series, the following steps are employed. First, the autoregressive and the moving average order, $r$ and $s$, respectively, are determined by minimizing model selection criteria (e.g. AIC). Next, the partial autocorrelations of the chosen ARMA($r,s$) process are parametrized in terms of the autoregressive and the moving average parameters, $\phi_{1},\ldots,\phi_{r},\theta_{1},\ldots,\theta_{s}$. Using the relation between autocorrelations and Kendall’s $\tau$, the Kendall’s partial autocorrelations are obtained. In a third step, all partial copulas in Eq. 5 are set to a chosen family (e.g. Gaussian or Gumbel) and are reparametrized by Kendall’s $\tau$ and consecutively by the AR and MA parameters, which can then be estimated by maximum likelihood. Bladt et al. (2025) generalize this approach and consider a Gaussian D-vine, i.e. all copulas in Eq. 5 are Gaussian, and then replace a finite amount of selected copulas by non-Gaussian copulas. Henceforth their approach will be called ’IPD’ for infinite partial dependence. The conceptual difference to our MAGMAR-copula model and to the model by Joe (2014) is that their approach is based on classical ARMA models. Furthermore, the partial copulas are modeled and not the AR- and MAG-copulas. We discuss the conceptual differences in more detail below.
@@ 2.4 Review of the extension by Joe
Now we want to briefly introduce the model of Joe (2014), which is to the best of our knowledge the earliest copula extension of the ARMA model. The model for a time series $\{U_{t}\}_{t\in\mathbb{Z}}$ with uniform marginals is given by the following updating equations:
@@
where $\varepsilon_{t}\stackrel{{\scriptstyle iid}}{{\sim}}U(0,1)$ are the common innovations, $h$ is the conditional quantile function of a $(p+1)$-variate copula and $g$ is the conditional quantile function of a $(q+1)$-variate copula. It is required that the copula corresponding to $h$ is translationally invariant and that the $q$-variate marginal distribution of the copula corresponding to $g$ is the independence copula. The idea is inspired by the Poisson ARMA process of McKenzie (1988): First create a process corresponding to an AR process, Eq. 9, and then feed this process to an MA-like filter together with the common innovations, as in Eq. 10. In our proposed MAGMAR-copula model, we too attempt to mix AR-type and MA-type copula processes. The difference, however, is that we feed the MA-type process (what we call ’MAG’) into the AR-type filter, similar to classical ARMA-type models .
2.5 Comparison of the Models with the Proposed Model
We briefly point out the main differences between the model by McNeil and Bladt (2022), the model by Joe (2014) and our proposed model. In A.7 we compare the models in detail w.r.t. flexibility, computational demand and modeling approach. For brevity we give the conclusions of the comparison here.
The MAGMAR and common-innovation model are similar in spirit. However, the MAGMAR-model has a more straight-forward modeling approach. In return, the common-innovations approach can be expected to be more tractable due to the stationary distribution naturally being standard-uniform. The additional adjustment transformation needed for the MAGMAR-model hinders theoretical tractability. The computational demand can be expected to be similar. The flexibility of the common-innovations approach is a potential direction for future research, since it is not obvious due to the AR-part being latent.
The modeling approach of the MAGMAR and IPD model are similar as both models can rely on e.g. autodependence and scatter plots. The main difference is that the partial copulas of $(U_{t},U_{t-k}|\mathcal{F}_{t-1:t-k+1})$ result from the AR and MAG-copula in the MAGMAR model, and are typically non-standard, whereas in the IPD model, the partial copulas can be chosen directly. This results in the MAGMAR model being preferable if the long-term dependencies are non-Gaussian. If the long-term dependencies are Gaussian and only a few partial dependencies seem to be non-linear, the IPD model is more suitable. While both models have moderate computational demand, the IPD models demand is less.
@@ 3 MAGMAR-Copulas
In this section, we introduce and discuss the moving aggregate modification to copula-based time series models. For illustration, the autoregressive and moving aggregate will be set to one. Results for arbitrary orders $p$ and $q$ can be derived in a similar fashion but are beyond the scope of this paper; however, some results with arbitrary orders $p$ and $q$ are given in the Appendix.
@@ 3.1 Introduction to the model
Our proposed approach enabling the model to overcome the Markov restriction amounts to modifying the autoregressive updating equation, Eq. 8, by a moving aggregate part. The moving aggregate part is, similar to classical, linear ARMA processes, a function of the current and past innovations. It is given as a conditional quantile function, corresponding to a copula, which we parametrize with parameter(s) $\theta$. So instead of $w_{t}$ in the updating equation, Eq. 8, we have $(R^{1}_{\theta_{1},\ldots,\theta_{q}})^{-1}((w_{t-q},\ldots,w_{t-1}),w_{t})$. Where the inversion is again w.r.t. the last element. We call the model Moving Aggregate Modified Autoregressive Copula-Based Time Series Model (MAGMAR-Copula model). Below we formally define the model with arbitrary model orders $p$ and $q$. The exposition that follows focuses on the case $p=q=1$. Furthermore, we assume that 1) all copulas are differentiable, 2) that they are within the Fréchet-Hoeffding bounds (see Joe, 2014, Chapter 2.9) and 3) that the inverses of $h$-functions associated to all used copulas exist. Now we formally introduce the model.
Definition 1 (MAGMAR(p,q)(p,q)-Copula Model).
@@
Let $R_{\phi_{1},\ldots,\phi_{p}}:(0,1)^{p+1}\rightarrow(0,1)$ be a Rosenblatt function associated to a stationary D-vine copula $C_{\phi_{1},\ldots,\phi_{p}}$, which is parametrized by parameters $\phi_{1},\ldots,\phi_{p}$. Furthermore, let $R^{-1}_{\theta_{1},\ldots,\theta_{q}}:(0,1)^{q+1}\rightarrow(0,1)$ be an inverse Rosenblatt function associated to a D-vine copula $C_{\theta_{1},\ldots,\theta_{q}}$ with independent margins, which is parametrized by parameters $\theta_{1},\ldots,\theta_{q}$. The time series $\{U_{t}\}_{t\in\mathbb{Z}}$ following the updating equation
@@
is a MAGMAR(p,q)(p,q)-time series. The corresponding model is the MAGMAR(p,q)(p,q)-Copula model. The copula $C_{\phi_{1},\ldots,\phi_{p}}$ is the AR-Copula. The copula $C_{\theta_{1},\ldots,\theta_{q}}$ is the MAG-Copula. The LHS of Eq. 11 is the AR-part and the RHS is the MAG-part. The time series following the updating equation
$U_{t}=(R^{1}_{\theta_{1},\ldots,\theta_{q}})^{-1}((w_{t-q},\ldots,w_{t-1}),w_{t})$ (12)
is the MAG(q)(q)-copula model. Furthermore, we define the adjusted MAGMAR(p,q)(p,q)-copula time series as $\tilde{U}_{t}=\Psi(U_{t}),\ \forall t\in\mathbb{Z}$, where $\Psi$ is a distribution function on $(0,1)$.
The Rosenblatt functions in Eq. 11 ought to be understood as a composition of $h$-functions (as given in Eq. 7). The AR-Rosenblatt function, $R^{1}_{\phi_{1},\ldots,\phi_{p}}$ is composed of $h$-functions $h_{\phi_{1}},\ldots,h_{\phi_{p}}$ while the MAG-Rosenblatt function is composed of $h$-functions $h_{\theta_{1}},\ldots,h_{\theta_{q}}$. The AR-copula is a stationary D-vine copula, cf. Sect. 2.2 or McNeil and Bladt (2022). The MAG-copula has independent margins. This property is also imposed by Joe (2014). The family of the copulas (e.g. normal, $t$, Gumbel, etc.) has to be chosen. Parameters $\phi$ and $\theta$ have to be set/estimated. We note that the MAG-part, that is used here, was also derived in McNeil and Bladt (2022) as the ’causal representation’ for the classical copula-based time series models. In Joe (2014), it is also given as the ’$q$ -dependent copula extension’. The new contribution is that we use this representation to construct the novel MAGMAR-Copula model. The reason for the adjustment, using a distribution function $\Psi$, is the following: The unconditional stationary distribution of a MAGMAR(p,q)(p,q)-copula time series, $\{U_{t}\}$ is in general not uniform if $p,q>0$. The distribution that arises is a distribution on $(0,1)$, and it depends on the AR and MAG parameters $(\phi,\theta)$. The probability integral transformation using $\Psi$ then ensures that $\tilde{U}_{t}\sim U(0,1)$. Notice that we did not prove that the stationary distribution of the MAGMAR-copula time series actually exists yet. The corresponding proposition is given in Sect. 3.3 for the MAGMAR(1,1)(1,1)-copula model. This is the reason for introducing $\Psi$ as a distribution function on $(0,1)$ and not as the stationary distribution of the process in Def. 1. In applications, we estimate $\Psi$ with non-parametric methods in a complementary step. In the following, we mostly consider the unadjusted time series $\{U_{t}\}$, since the transformation does not affect most results. When appropriate, we also give the results for the adjusted time series.
Before giving the relations to other models and deriving distributional properties, we proceed to discuss implications of the model equation and further explain the model for the case where $p=q=1$. In this case the updating equation is given as
$U_{t}=h_{\phi}^{-1}(h_{\theta}^{-1}(w_{t},w_{t-1}),U_{t-1}).$ (13)
@@
First, we note that $C_{\theta}$ is not the copula of $w_{t}$ and $w_{t-1}$. The innovations $w_{t}$ and $w_{t-1}$ are independent and their copula is the independence copula, $F_{w_{t},w_{t-1}}=\Pi$. The expression $h_{\theta}(w_{t},w_{t-1})$ is to be understood as a function of $(q+1)$ random variables $w_{t}$ and $w_{t-1}$, just as $\varepsilon_{t}-\theta\varepsilon_{t-1}$ in the classical ARMA model is a function of two random variables $\varepsilon_{t}$ and $\varepsilon_{t-1}$. Second, the copula $C_{\phi}$ is also not the copula of $U_{t}$ and $U_{t-1}$, however it is connected to the copula of the two random variables. This is analogous to the classical ARMA(1,1) model, where the partial autocorrelation of $X_{t-1}$ and $X_{t}$ is not solely determined by the AR(1) parameter but also by the MA(1) parameter. (The influence of the MAG-Copula and the AR-Copula to the dependence of $U_{t}$ and $U_{t-1}$ becomes clear when we derive the joint conditional distribution of $U_{t}$ and $U_{t-1}$ given $\mathcal{F}_{t-2}$.) The expression $h_{\phi}(h_{\theta}^{-1}(w_{t},w_{t-1}),U_{t-1})$ may also be understood as a function of random variables. To verify that the MAGMAR(1,1)(1,1) model equation actually yields the expected behavior to have no Markov restriction i.e. an implicit infinite autoregressive order, the model equation can be iterated. To this aim, an explicit form for the innovations is necessary. It can be directly obtained from Eq. 13 as
@@
Hence, in general, each past observation, $U_{t-i},i=1,2,\ldots$ has an actual influence on $U_{t}$. This is the desired property. Furthermore, the MAG($\infty$) representation can be derived as
@@
As noted before, there also exists a MAG($\infty$) representation for the case with no moving aggregate extension. The representation was derived in McNeil and Bladt (2022) as ’causal representation’. It is easy to see that for $C_{\theta}=\Pi$, the two representations are equal. This is in line with proposition 1, which proposes that for $C_{\theta}=\Pi$, the MAGMAR(1,1)-Copula model is equal to the classical copula-based time series model. The corresponding proposition is given in Sect. 3.2. In Sect. 3.3, it is explicitly shown that the MAG($\infty$) representation of the MAGMAR-Copula time series converges for the case where the AR-copula, $C_{\phi}$, is the normal copula. For other copulas, it is hard to establish the convergence of the MAG()($\infty$) representation explicitly. However, we can use results from Douc et al. (2014), involving a contraction-on-average, to establish desired properties.